Evaluate the determinant in the matrix
|4 8|
|5 6|

Answer by Alan3354(69443)

Do it like this one:
----
Evaluate the determinate in the matrix.
|1 -3|
|-8 2|
----------
Det = 1*2 - (-8*-3) = 2 - 24
= -22